### v1

# Guarding the Guardrails: A Taxonomy-Driven Approach to Jailbreak Detection

Abstract
Jailbreaking techniques pose a significant threat to the safety of Large Language Models (LLMs). Existing defenses typically focus on single-turn attacks, lack coverage across languages, and rely on limited taxonomies that either fail to capture the full diversity of attack strategies or emphasize risk categories rather than the jailbreaking techniques. To advance the understanding of the effectiveness of jailbreaking techniques, we conducted a structured red-teaming challenge. The outcome of our experiments are manifold. First, we developed a comprehensive hierarchical taxonomy of 50 jailbreak strategies, consolidating and extending prior classifications into seven broad families, including impersonation, persuasion, privilege escalation, cognitive overload, obfuscation, goal conflict, and data poisoning. Second, we analyzed the data collected from the challenge to examine the prevalence and success rates of different attack types, providing insights into how specific jailbreak strategies exploit model vulnerabilities and induce misalignment. Third, we benchmark a popular LLM for jailbreak detection, evaluating the benefits of taxonomy-guided prompting for improving automatic detection. Finally, we compiled a new Italian dataset of 1364 multi-turn adversarial dialogues, annotated with our taxonomy, enabling the study of interactions where adversarial intent emerges gradually and succeeds in bypassing traditional safeguards.
1 Introduction
Large Language Models (LLMs) often exhibit unintended behaviors such as hallucinations, biased or toxic outputs, or even responses that may compromise the security of the system in which the model is deployed. These behaviors represent instances of misalignment, which refers to a deviation from the intended objective of being both helpful and safe. Preventing misalignment is critical for LLMs that are integrated into real-world applications, and it remains a central concern in safety research.
Despite efforts to align LLMs with human preferences through Supervised Fine-Tuning (SFT), often followed by Reinforcement Learning from Human Feedback (RLHF) Ziegler et al. (2020); Stiennon et al. (2020); Ouyang et al. (2022) or Direct Preference Optimization (DPO) Rafailov et al. (2024), these models can still generate unsafe content, even in response to benign user inputs. As shown by Betley et al. (2025), even small perturbations in fine-tuning, such as a single epoch of training on insecure code, can lead to significant misalignment. These risks are further amplified by adversarial attacks, where malicious actors exploit the model’s vulnerabilities to induce harmful outputs.
One major challenge in ensuring model safety is the phenomenon of jailbreaking, a form of adversarial prompting in which the model is manipulated into misalignment. While some jailbreaks focus on crafting a single malicious prompt, others unfold over the course of several turns. These multi-turn jailbreaks Russinovich et al. (2025) gradually steer the model toward the desired outcome through a series of benign-looking steps, making detection particularly difficult since the malicious intent is distributed across the interaction.
To mitigate the risks of misalignment and jailbreaking, guardrailing systems are built as safety layers around the core language model. These systems monitor, constrain, or intervene in the model’s behavior to prevent undesired outputs. Common components include anomaly detectors, prompt sanitizers, decoding constraints, and other filters Jain et al. (2023); Cao et al. (2024); Zeng et al. (2024). Among these, external safety modules play a central role. Examples include content moderation tools, such as the OpenAI Content Moderation API, [^1] https://platform.openai.com/docs/guides/moderation/overview Perspective API, [^2] https://perspectiveapi.com/ and Llama Guard Inan et al. (2023). These detectors are typically implemented as trained classifiers or specialized LLMs fine-tuned on safety-related data to recognize and block malicious activity before harm occurs.
While guardrailing systems provide essential protective layers, their reliability depends heavily on the robustness of their individual components. In particular, adversarial attack detectors must be thoroughly examined before being integrated into safety infrastructures. The more accurate and generalizable they are, the better they can proactively identify emerging jailbreak attempts. To be effective, such systems must cover a broad spectrum of attack strategies across domains and languages. Existing defenses lack robustness against multi-turn jailbreaks Li et al. (2024), as they are assessed only on single-turn adversarial prompts, which represents a threat model that fails to reflect real-world dynamics. Training detectors capable of handling multilingual and multi-turn attacks requires curated datasets with annotated adversarial prompting strategies grounded in a comprehensive taxonomy. However, such data are scarce or unavailable for most languages, including Italian.
Moreover, existing taxonomies primarily emphasize the type of harm produced by an attack rather than the prompting technique that generates it. Others are narrowly focused on a single class of attacks, such as persuasion Zeng et al. (2024), and therefore fail to capture the full diversity of jailbreak strategies. In addition, jailbreak approaches evolve rapidly, and taxonomies that are not updated accordingly quickly lose their relevance. These limitations reduce the usefulness of existing taxonomies for annotation, detector training, and targeted mitigation.
To address these gaps, we present an Italian dataset of 1364 unsafe multi-turn dialogues spanning a wide range of jailbreaking techniques, collected through a structured red-teaming challenge. We propose a hierarchical taxonomy comprising 50 jailbreaks that overcomes the limitations of existing taxonomies and use it to annotate the dataset. Leveraging this resource, we analyze the prevalence of different attack types and evaluate the ability of GPT-5 to detect these adversarial behaviors.
Our overall objective is to contribute to the enhancement of guardrailing systems in which an attack detector plays a central role. We present four main contributions:
  1. 1.
    We release a new dataset for evaluating the safety and performance of adversarial prompt detectors in Italian. The dataset covers both single-turn and multi-turn jailbreaks and addresses the critical scarcity of such resources in the field. To the best of our knowledge, this is the first dataset that is simultaneously in Italian, multi-turn, and annotated specifically for jailbreak detection. The dataset and the material will be releases upon acceptance.
  2. 2.
    We propose a comprehensive taxonomy of jailbreaking techniques against LLMs. This taxonomy aggregates and extends existing categorizations from the literature and is further refined through empirical observations of real attacks collected during dataset construction, resulting in broader and more detailed coverage than previous approaches.
  3. 3.
    We share insights from our analysis of the data collected using the proposed taxonomy, including success rates of different techniques, and the impact of combining them.
  4. 4.
    We evaluate the comparative performance of GPT-5 in adversarial attack detection with and without taxonomy-enhanced prompting across two complementary settings. This evaluation establishes a structured methodology for testing adversarial attack detectors and provides empirical evidence on the benefits of integrating a taxonomy into the prompting process.
The remainder of this paper is structured as follows. In Section 2, we review the related work. Section 3 outlined our red teaming challenge for dataset construction (3.1) and the taxonomy design (3.2). The results are presented in Section 4, followed by the findings from our use-case experiments in Section 5. Finally, Section 6 concludes the paper and discusses future research directions.
2 Related work
Jailbreak datasets
To the extent of our knowledge, no existing dataset combines Italian language, multi-turn dialogues, and explicit jailbreak-type annotations. Deng et al. (2024) introduce MultiJail, a multilingual dataset constructed by manually translating English jailbreak prompts into nine languages, including Italian, while Pernisi et al. (2024) explore jailbreaking in Italian through a many-shot prompting technique, an extension of few-shot prompting that includes numerous demonstrations of unsafe behavior within a single prompt. However, both datasets are limited to single-turn interactions and are annotated for harm categories rather than adversarial strategies.
In English, several datasets provide multi-turn conversations labeled for safety. Some are synthetic and annotated primarily for harm rather than adversarial strategy. For example, CoSafe by Yu et al. (2024a) consists of GPT‑4-generated dialogues simulating coreference-based attacks, labeled with binary harmfulness judgments. Ung et al. (2022) collect real human-model conversations and annotate them to capture evolving safety dynamics, but without categorization of attack types. While some datasets do annotate for adversarial strategies, such labels are not always publicly released. Ganguli et al. (2022), for example, present the AnthropicRedTeam dataset, which consists of human-generated red teaming transcripts and features rich internal annotations, including tags describing the adversarial techniques. However, these labels remain inaccessible to the research community.
Public datasets that explicitly label adversarial techniques in multi-turn dialogue are relatively rare. SafeDialBench Cao et al. (2025) and Multi-Turn Human Jailbreaks (MHJ) Li et al. (2024) offer public multi-turn datasets annotated with 7 distinct jailbreaking techniques each. In MHJ, these labels are informed by red teamers’ own metadata describing their rationale and strategy. Both resources contribute into adversarial prompting in multi-turn setting, yet their restricted taxonomies and focus on English (with Chinese in SafeDialBench) underscore the continued lack of wider coverage.
Jailbreaking taxonomies
A variety of taxonomies address different aspects of LLM safety. Many efforts classify jailbreaks by the type of risk they exhibit Rao et al. (2024); Weidinger et al. (2022); Geiping et al. (2024). As LLMs become increasingly embedded into downstream applications, several works concentrate on the infrastructural risks they entail, introducing taxonomies of indirect prompt injection in integrated systems Greshake et al. (2023) and frameworks proposing unified classifications that cover both model-level and infrastructure-level attacks Zahid et al. (2025).
Across methodological approaches, a key distinction is often drawn between automatic and human-crafted jailbreaks. Yi et al. (2024) classify attacks by setting (black-box versus white-box), and by automation level (manual versus optimized). Similarly, Chu et al. (2025) categorize by construction method, including human designed prompts and optimization driven ones. In this work, we adopt the black-box setting and focus primarily on interpretable, human-crafted prompts. Unlike the above approaches, our taxonomy emphasizes the linguistic and strategic mechanisms through which jailbreaks succeed, rather than the type of risk or attack surface involved. Existing taxonomies in this line of research range from narrowly focused to more general frameworks.
Narrow-focused studies analyze specific jailbreak strategies in depth. For instance, Wei et al. (2023) identify two key alignment failure modes exploited by jailbreaks: competing objectives and mismatched generalization, while Zeng et al. (2024) examine persuasion mechanisms as a targeted attack vector. In contrast, broader generalization efforts aim to systematize a wider variety of techniques. HackAPrompt Schulhoff et al. (2023) stands out as a key foundation, collecting 600K adversarial prompts through a jailbreaking competition. In a similar direction, Liu et al. (2024) propose a taxonomy of ten techniques organized into three families. Yu et al. (2024b) categorize jailbreak prompts into five categories and ten patterns, grounding their analysis in strategies observed among real users. Finally, Rao et al. (2024) link specific prompting techniques to the underlying adversarial intent. While these taxonomies provide valuable perspectives, none achieve full coverage of the diverse jailbreak strategies observed in practice.
3 Red teaming challenge and taxonomy
Our methodology builds on the red teaming paradigm, which typically relies on human experts or LLMs to probe for unsafe behaviors, thereby exposing potential vulnerabilities and biases and informing system improvements. In our work, we combine a structured human red teaming challenge with a taxonomy-based annotation framework. This leads to the creation of a multi-turn adversarial dialogue dataset, with each conversational thread annotated with the attack techniques it contains.
3.1 Red teaming challenge
We organized a structured red teaming challenge involving 48 participants from the Master’s course on Seminars on AI and Robotics at Sapienza University of Rome. Each participant had a two-hour session to perform multi-turn adversarial attacks. The target model was Minerva-7B-instruct-v1.0, [^3] Minerva is a family of LLMs developed by Sapienza NLP in the context of the Future Artificial Intelligence Research (FAIR) project, in collaboration with CINECA. an instruction-tuned LLM pretrained on Italian and English corpora.
Participants were divided into groups and assigned tasks corresponding to one of three vulnerability areas: Attacks on Data, Attacks on the Model, and Attacks on Infrastructure. Across these areas, nine tasks were defined. Attacks on Data included eliciting gender bias, eliciting ethnicity bias, and inducing privacy violations. Attacks on the Model comprised generating information that could cause physical or non-physical harm and triggering hallucinations. Finally, Attacks on Infrastructure involved bypassing arbitrary restrictions defined by the system prompt, revealing a hidden word, and extracting fragments of the prompt itself. To support the infrastructure-related tasks, the system prompt was modified with explicit prohibitions, such as arbitrary restrictions, requiring that the hidden word remains undisclosed, and forbidding any disclosure of the prompt itself.
The attacks were conducted primarily in Italian, reflecting the optimization of the model, with a small portion in English. In total, 1364 adversarial conversations were collected and annotated. These conversations are multi-turn, containing an average of 2.95 adversarial prompts each. We manually annotated the conversations using the proposed taxonomy, grounded in the literature and refined with observations from the challenge. The annotation was designed to capture combinations of techniques.
3.2 Taxonomy
We present a comprehensive taxonomy of 50 prompt-based jailbreaking techniques targeting Large Language Models, illustrated in Figure 1. This taxonomy consolidates and extends prior classifications from the literature, integrating insights from existing taxonomies and further refining them through empirical observations of attacks collected during our red-teaming challenge.
The taxonomy is organized into three hierarchical levels and groups techniques into seven distinct families, each defined by the primary mechanism through which adversarial prompts bypass safety safeguards: Impersonation Attacks & Fictional Scenarios, Privilege Escalation, Persuasion, Cognitive Overload & Attention Misalignment, Encoding & Obfuscation, Goal-Conflicting Attacks, and Data Poisoning Attacks. In the following subsections, we discuss each family in detail, outlining its underlying mechanisms and representative jailbreak strategies.
Figure 1: Visual representation of the proposed hierarchical taxonomy of prompt-based jailbreak techniques against LLMs, showing seven mechanism-driven families and their subcategories.
3.2.1 Impersonation Attacks & Fictional Scenarios
This family of attacks induces the model to assume roles or operate within fictional contexts that relax its alignment constraints. This pattern is widely studied in the literature and corresponds to categories found in several taxonomies, including Cognitive Hacking (COG) Rao et al. (2024), Pretending family Liu et al. (2024), and Virtualization category Kang et al. (2023). Instances include:
  • Role Play: The model is prompted to act as specific individuals (e.g., a malicious expert, a criminal, or an unfiltered AI), creating implicit associations between certain roles and unmoderated behavior. Further subdivisions are identified by Yu et al. (2024b), which distinguish between Defined Personas and Virtual AI, with the latter further split into Superior, Opposite, and Alternate modes.
  • Benign Context Framing: Malicious requests are embedded within carefully crafted scenarios that appear to carry lower perceived risk. This includes, for instance, presenting the request as part of academic research, sociological analysis, or a controlled experiment. This category corresponds to the Disguised Intent patterns described in Yu et al. (2024b). A specific variant frames the request within a Game scenario, which in our taxonomy is treated as a third-level subcategory.
  • Fictional Framing: Harmful requests are presented within jokes, stories, or imagined scenarios, making them appear legitimate and creative. This category maps to Imagined Scenario and partially overlaps with Joking Pretext in Yu et al. (2024b).
These techniques are often combined to create more sophisticated jailbreaks. Role Play, in particular, is among the most used approaches and constitutes the basis of several prominent prompt families.
3.2.2 Privilege Escalation
Privilege escalation attacks simulate elevated privileges or unconstrained execution contexts to induce the model to bypass its safety restrictions. Typical techniques include claiming administrative roles, declaring that the model has been jailbroken, or using formatting cues that reinforce perceived command authority. This class of attacks corresponds to the one in Liu et al. (2024) and includes:
  • Sudo/Admin Mode: The prompt asserts that the model is running in a privileged mode (e.g., “developer”, or “sudo”), implying that it should respond without constraints. A variation consists of masking the request behind a “special” instruction Schulhoff et al. (2023).
  • Jailbroken Model Simulation: The model is explicitly told that it has been jailbroken or freed from its constraints and should therefore comply with otherwise restricted requests.
  • Typographical Authority Simulation: Requests are written in uppercase or include other visual cues that simulate authority. Although simple, such signals have been empirically observed to increase compliance by mimicking the style of commands or urgent directives.
3.2.3 Persuasion
Large Language Models can be induced to produce unsafe outputs through the use of persuasive language. Trained on extensive corpora of human dialogue, they implicitly acquire patterns of social influence and negotiation, which can be exploited to bypass alignment safeguards.
Zeng et al. (2024) present a comprehensive analysis of adversarial persuasion, identifying forty techniques grouped into thirteen strategies. Building on this framework and including observations from other works, we distill the strategies most directly relevant to jailbreaks into six primary second-level categories within the Persuasion family.
  • Logical, Evidential, and Quantification-Based Persuasion: Prompts leverage logic or quantitative data to achieve malicious goals. By presenting requests as rational or evidence-based, they exploit the model’s tendency to comply with seemingly factual reasoning.
  • Authority and Norm-Based Persuasion: Prompts invoke real or fabricated authority, citing trusted sources such as domain experts to legitimize unsafe requests Yang et al. (2024).
  • Emotional, Reciprocity-Based, and Commitment-Based Persuasion: These techniques mimic interpersonal dynamics between the user and the model, leveraging emotions, praise, or references to past cooperation. They often suggest a social obligation to comply, inducing feelings of reciprocity or debt. A common variation is the Repeated Request technique, where the attacker asserts that the model has previously fulfilled the same request.
  • Instruction Repetition: The attacker repeats the same instruction multiple times, appearing as “insisting” until the model complies Rao et al. (2024). This approach can make the request appear more acceptable and has been studied as a persuasion dynamic.
  • Urgency and Scarcity-Based Persuasion: Harmful requests simulate urgency or limited resource availability, creating artificial pressure that increases the likelihood of compliance.
  • Manipulative and Coercive Persuasion: The most overtly adversarial form of persuasion, pressuring the model into unsafe behavior using coercion or invoking negative consequences.
3.2.4 Cognitive Overload & Attention Misalignment
These attacks bypass moderation by creating complex or overwhelming contexts that divert the attention of the model away from safety constraints. They exploit both computational and attentional limitations. This class corresponds to the Attention Shifting category described by Yu et al. (2024b).
  • Distractor Instructions: Innocuous and deceptive objectives are combined to mislead the model. This category maps to the Distractor/Negated Distractor defined in Wei et al. (2023).
  • Mathematical & Decomposition Attacks: Malicious requests are reformulated as mathematics or multi-step logical problems Bethany et al. (2024), or decomposed into fragments that the model is later asked to recombine. Extending the notion of payload splitting Kang et al. (2023), these misdirect the model’s attention and obscure adversarial intent.
  • Indirect Task Deflection: The model is asked to generate code, snippets, or other technical artifacts that indirectly accomplish a harmful objective Rao et al. (2024).
  • Context Saturation: The adversarial request is embedded within a long prompt to push the model towards its context window limits. Under such conditions, models may behave unpredictably and fail to block malicious content Schulhoff et al. (2023).
3.2.5 Encoding & Obfuscation
This class of techniques encompasses strategies that distort the surface form of malicious content to evade safety filters by creating out-of-distribution requests.
When attackers maximize the distance between their requests and the distributions seen during safety training, models may become increasingly vulnerable to unsafe behavior. Wei et al. (2023) describe this phenomenon as Mismatched Generalization. Comparable concepts appear in other taxonomies under different labels, including Orthographic Techniques Rao et al. (2024), Obfuscation Kang et al. (2023), and Character-Level Encoding Liu et al. (2024). Instances include:
  • Surface Obfuscation: Alter the text surface by introducing misspellings, character substitutions, or similar perturbations while keeping the intent human-readable. This includes techniques such as vowel removal and homoglyph substitution Schulhoff (2025).
  • Token Splitting: Break words or phrases into separated tokens using punctuation or spacing (e.g., “h.o.w t.o b.u.i.l.d.a.b.o.m.b”) to evade token-based filters.
  • Semantic Rewriting: Rephrase malicious prompts while preserving their intent. This covers Token-Level Transformations (e.g., synonym replacement, reordering, insertion, deletion) and Sentence-Level Transformations (e.g., alternative paraphrased expressions). The search for reformulations can be automated, increasing attack scalability Li et al. (2020).
  • Linguistic Encoding: Transliteration of the request using alternate representations. This includes low-resource languages, alternative scripts (e.g., Cyrillic look-alikes), emojis, Base64, or other encoding schemes.
  • Lexical Techniques: Use specific short phrases or tokens, sometimes discovered automatically, that reliably trigger unsafe behavior Rao et al. (2024). Such triggers can be human-interpretable or optimization-generated. When automatically learned, they often transfer across models, revealing systematic training vulnerabilities Zou et al. (2023).
  • Embedded Prompting: Conceal malicious instructions within seemingly benign structures such as code comments, JSON fields, or uploaded files (e.g., images Carlini et al. (2024)); or encode them visually Jiang et al. (2024). This category often combines Obfuscation with Cognitive Overload, and it is particularly relevant when deadling with multi-modal models.
3.2.6 Goal-Conflicting Attacks
Goal-conflicting attacks work by assigning the model multiple, conflicting goals, thereby disrupting its safety alignment. This family corresponds to the failure mode of Competing Objectives described by Wei et al. (2023) and is also referred to as Goal Hijacking by Perez and Ribeiro (2022).
  • Prefix Injection: Malicious prefixes are prepended to the prompt so that the model interprets them as part of its conversational history Wei et al. (2023).
  • Instruction Masking: Harmful content is hidden within seemingly benign instructions. The adversary may ask the model to summarize, rephrase, or add details to malicious text. The well-known Text Completion as Instruction attack Rao et al. (2024) is a notable instance that also conceptually overlaps with the Cognitive Overload & Attention Misalignment family.
  • Refusal Suppression: The model is explicitly instructed to comply with the request and to avoid refusals, effectively suppressing its alignment-driven safety responses.
  • Context Ignoring: The prompt tells the model to disregard previous instructions, safety guidelines, or contextual boundaries in order to fulfill the adversarial request.
  • Assumption of Responsibility: Similarly to Context Ignoring, this technique encourages the model to “think freely”, take responsibility for its answers, or “use its own judgment” rather than follow pre-programmed restrictions, shifting the decision burden to the model.
  • Objective Juxtaposition: The prompt combines legitimate objectives with harmful ones, creating an internal goal conflict. This paring can override safety.
3.2.7 Data Poisoning Attacks
Data Poisoning Attacks aim to corrupt the behavior of the model by manipulating its conversational context. Instead of directly issuing an explicit harmful request, these techniques guide the model toward unsafe outputs by introducing unaligned examples, false premises, or gradually escalating elements that can later push it to produce harmful content.
  • Incremental Poisoning: The malicious request is distributed across multiple turn, progressively increasing in harmfulness, often starting with innocent prompts.
  • Many-Shot Jailbreaking: Exploits in-context learning by providing numerous adversarial prompt-response pairs in which the model complies with harmful requests, thus inducing unaligned behavior Anil et al. (2024); Pernisi et al. (2024).
  • False Fact/Bias Instillation: Injects fabricated information or biased premises into the conversational context.
4 Challenge data analysis
Our dataset, collected during the Red Teaming challenge, consists of 1364 unique branches of adversarial conversations, of which 185 correspond to successful attacks. On average, each user-assistant interaction contains 2.85 messages and 176.51 words. For successful cases, these averages slightly change to 3.02 messages and 168.57 words per conversation. This corpus allows us to derive insights of the effectiveness of different adversarial prompting strategies, analyzed here by level.
Figure 2: Distribution of adversarial dialogues across first-level jailbreak families, showing total occurrences and successful attacks for each category.
4.1 First-level jailbreak techniques
Figure 2 illustrates the distribution of adversarial dialogues across the first-level jailbreak categories, showing both the number of occurrences and corresponding successful cases. Detailed results, including success rates, are reported in Table 1.
The most prevalent jailbreak family employed during the red-teaming challenge was Impersonation Attacks & Fictional Scenarios, which appeared in 696 dialogues (51.0% of the total). The Data Poisoning Attacks family achieved the highest success rate (17.2%), while Encoding & Obfuscation techniques showed the lowest (9.4%), having minimal effect on the tested model except for the Lexical Techniques subcategory. Figure 2 and Table 1 also include two auxiliary categories: No Technique and Automated Attacks. The No Technique category accounts for cases in which participants successfully completed jailbreak tasks without applying any explicit attack strategy, directly issuing request to the model. Including this category highlights how the use of targeted jailbreak techniques significantly increases the overall success rate of adversarial attempts.
Finally, the Automated Attacks category represents an orthogonal dimension relative to our taxonomy. In our red teaming challenge, participants could rely only on adversarial prompting and not on optimization-based methods, given the limited time and resources available. However, automatically discovered triggers, previously identified in other studies as transferable across models Zou et al. (2023), were permitted for testing. These Automatic Attacks achieved the highest success rate among all categories, as reported in Table 1.
4.2 Second-level jailbreak techniques
Most interactions in the dataset are annotated with multiple labels, reflecting that jailbreaks often rely on combining complementary techniques to maximize their effectiveness. For this reason, our analysis examines both isolated and combined uses of techniques.
In isolated use, Benign Context Framing is the most frequent second-level category (51 occurrences), followed by Lexical Techniques attack (41) and Incremental Poisoning (36). Benign Context Framing was also used by the largest number of users (36 unique participants), and is the only technique present in at least one successful attack for each of the nine challenge tasks. The techniques with the highest number of successful attacks were Lexical Techniques and Incremental Poisoning (12 each), emphasizing the potency of complex multi-turn strategies.
When examining second-level techniques individually, but as components of combined attacks rather than isolated ones, Role Play emerges as the most frequent technique, occurring 331 times, of which 240 instances belong to the Virtual AI variant. It is followed by Context Ignoring (244 occurrences) and Benign Context Framing (240 occurrences). Prefix Injection stands out with a success rate of 31.1% (19 successful dialogues), followed by Objective Juxtaposition with 13 successful cases. Together, they form the most effective multi-technique pair, with 6 successes out of 20 conversations.
Several predefined multi-technique jailbreak prompts also demonstrated notable effectiveness. The Khajiit [^4] ChatGPT Khajiit Jailbreak Prompt group and DAN (“Do Anything Now”) family Shen et al. (2024) showed particularly strong results. DAN prompts, which combine Fictional Framing with elements of Goal-Conflicting Attacks, appeared 22 times and succeeded in 7 cases across all tasks. Excluding the physical harm promotion task, the DAN approach achieved the highest success rate overall (31.8%). Across the four most common adversarial evaluation (physical and non-physical harm promotion, secret world disclosure, and system prompt extraction), DAN emerges as the most successful composite in absolute terms.
Table 1: Distribution of jailbreak families across all and successful conversations, with corresponding Success Rates (SR) for each category.
Jailbreak Family Conversations Successful Attacks SR (%)
Automatic Attacks 51 12 23.5
Data Poisoning Attacks 198 34 17.2
Impersonation Attacks & Fictional Scenarios 696 105 15.1
Persuasion 250 36 14.4
Cognitive Overload & Attention Misalignment 289 38 13.1
Goal-Conflict Attacks 624 74 11.9
Privilege Escalation 100 11 11.0
Encoding & Obfuscation 139 13 9.4
No Technique 110 7 6.4
5 Use case experiments
Finally, we present preliminary experiments aimed at exploring the potential benefits of using our taxonomy for improving adversarial attack detection. Specifically, we design two use case studies: Jailbreaking Attempt Detection and Jailbreaking Techniques Detection. In the first, the model must determine whether a user is attempting to jailbreak the system. In the second, the model must identify the jailbreaking techniques used. For these experiments, we employ GPT-5 as the backbone detector.
The core idea is to measure whether providing the model with our taxonomy improves its ability to identify jailbreaks. Beyond its use in dataset annotation, a taxonomy can also guide model behavior when included in the system prompt during training or evaluation Inan et al. (2023). While our dataset is not large enough to train a jailbreak detector analogous to Llama Guard, we aim to gain initial evidence of the potential impact of taxonomy-enhanced prompting for this purpose.
Both tasks are evaluated using our annotated dataset. We select only the dialogues where the jailbreak was successful and exclude interactions corresponding to infrastructural attacks, since in these cases the estimation of malicious intent is subjective. After this filtering, we obtain 91 records.
5.1 Jailbreaking attempt detection
In this experiment, GPT-5 is given the user turns from a user-assistant interaction and must determine whether the user is attempting to jailbreak the model. The model is instructed to name the jailbreaking technique it detects if it believes the user is attempting a jailbreak, or to return “benign” otherwise.
We first evaluate the effect of taxonomy enhancement using the transition matrix shown in Table 2. Here, a “transition” reflects how the model’s judgment about whether a jailbreak attempt is present changes once it is provided with the taxonomy. As shown in the upper-right cell, in 14.3% of cases the detector correctly identified a jailbreak attempt only when given the taxonomy. Conversely, in two instances performance decreased. Upon inspection, these appear to be reasonable misclassifications. Overall, the improvement is evident: the detection success rate increased from 65.9% without the taxonomy (left column) to 78.0% with taxonomy guidance (top row).
Table 2: GPT-5 jailbreak attempt detection results with and without taxonomy-enhanced prompting. Each cell reports the number and percentage of instances transitioning between detection outcomes.
Detected w/o taxonomy Undetected w/o taxonomy
Detected w/ taxonomy 58 (63.7%) 13 (14.3%)
Undetected w/ taxonomy 2 (2.2%) 18 (19.8%)
Figure 3: Jailbreaking attempt detection success rates by task w/ and w/o taxonomy enhancement.
To further investigate whether the improvement depends on the jailbreaking objective, we compare success rates by task before and after taxonomy enhancement, as illustrated in Figure 3. The privacy violation task is excluded due to insufficient data. For the remaining tasks, success rates increase consistently, with the largest gain (29.4%) observed in hallucination-inducing attacks. The only exception is the non-physical harm task, where the difference is minimal.
5.2 Jailbreaking techniques detection
This time, the model is provided with the full sequence of user-assistant turns, excluding only the assistant’s final response. Its task is to recognize successfully applied jailbreaking techniques that are likely to cause the assistant to comply with a restricted request in the next turn.
Because this task is inherently multi-class and multi-label, we need a systematic way to evaluate the detector’s free-form outputs before and after taxonomy enhancement. To enable quantitative comparison, we map the free-text labels generated by GPT-5 (when not provided with the taxonomy) to the closest categories in our taxonomy. We report recall as the primary metric for this experiment, as it reflects how many of the ground truth labels were correctly identified. For an adversarial attack detector, high recall is crucial: in a decision-making system that relies on the output of the detector, low recall implies that malicious requests could slip through undetected. By contrast, low precision, while undesirable, poses a less severe risk, as it merely results in benign prompts being unnecessarily blocked. As demonstrated in Table 3, the recall of GPT-5 consistently improves across all taxonomy levels after alignment with our taxonomy.
Table 3: Average Recall of GPT-5 in the Jailbreaking Techniques Detection task without and with taxonomy-enhanced prompting, reported across three hierarchical levels of the taxonomy.
Prompting Avg. Recall: lvl 1 Avg. Recall: lvl 2 Avg. Recall: lvl 3
Baseline 0.22 0.14 0.17
Taxonomy-enhanced 0.26 0.20 0.23
6 Conclusion and future works
This work provides new insights into the mechanisms and dynamics of multi-turn jailbreaking attacks, highlighting their incremental nature and the effectiveness of specific technique combinations. We introduced a comprehensive hierarchical taxonomy that achieves the broadest coverage of jailbreak strategies to date and applied it in constructing the first Italian dataset of multi-turn adversarial dialogues. Together, these contributions form a reproducible framework for studying adversarial prompting in safety-critical settings. Beyond its descriptive value, the proposed taxonomy demonstrated practical utility in improving the performance of adversarial attack detectors, a key component of modern guardrailing systems that safeguard large language models.
Looking forward, we plan to deepen the analysis of the incremental and temporal aspects of multi-turn attacks.To support this goal, a second edition of the red teaming challenge is planned, aimed at collecting longer dialogue trajectories and incorporating the currently underrepresented family of automated attacks. We also intend to maintain and expand the taxonomy as new jailbreak techniques emerge, ensuring that it remains a relevant and useful resource for the research community.